OpenStudy (anonymous):

Suppose you are trying to find the height of a triangle given the area and the length of the base. Rewrite the area formula below to solve for the height of the triangle a=1/2 bh

OpenStudy (anonymous):

We can assume that the base is greater than 0, otherwise we wouldn't have a triangle. Thus we can divide by 1/2 base without worrying about division by 0. \[a = \frac{1}{2}bh \Rightarrow \frac{a}{\frac{1}{2}b} = h \Rightarrow h = 2\frac{a}{b}\]